You borrow $456,996.03 at an annual interest rate of 4.12%, compounded once every year. If you have 5 years to pay off the loan, what must be your per period payment?
Answer #1

Period Payment = Loan / PVAF(r%, n)

PVAF is sum of PVF(r%,n)

PVAF = PVF (4.12%,1) + PVF(4.12%,2) +... + PVF(4.12%,5)

PVF(4.12%, 1) = 1 / (1.0412^1)

PVF(r%,n) = 1 / ( 1 + r)^n

= $ 456,996.06 / PVAF (4.12%, 5)

= 456,996.06 / 4.4369

= $ 103,000

Annual Paymnet is $ 103,000